Suppose the force acting on a column that helps to support a building is a normally distributed random variable X with mean value 14.0 kips and standard deviation 1.50 kips.

Compute the following probabilities by standardizing and then using a standard normal curve table from the Appendix Tables. (Round your answers to four decimal places.).

(b) P(X ≤ 17)

(c) P(X ≥ 6.5)

(d) P(13 ≤ X ≤ 16)

(e) P(|X − 14| ≤ 2)
Answer #1

a)

for normal distribution z score =(X-μ)/σx
here mean=       μ= 14
std deviation   =σ= 1.5000
probability = P(X<17) = P(Z<2)= 0.9772

b)

probability = P(X>6.5) = P(Z>-5)= 1-P(Z<-5)= 1-0= 1.0000

c)

probability = P(13<X<16) = P(-0.67<Z<1.33)= 0.9082-0.2514= 0.6568

d)

P(|X-14|<2) =P(-2+14 <X<2+14)=P(12<X<16)=P(-1.33 <Z<1.33)=0.9082-0.0918 =0.8164